40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Tue Oct 27 23:47:53 2020, gautam, Update, General, ISS checkout 
    Reply  Wed Oct 28 14:01:03 2020, not gautam, Update, General, ISS checkout 
       Reply  Wed Oct 28 14:07:47 2020, gautam, Update, General, ISS checkout 
Message ID: 15645     Entry time: Tue Oct 27 23:47:53 2020     Reply to this: 15647
Author: gautam 
Type: Update 
Category: General 
Subject: ISS checkout 

I wanted to look into the ISS situation. Some weeks ago, I found the PD that was previously used as the in-loop photodiode. I wanted to use this and measure the open-loop RIN at a few places (to see if there's any variation and also to check its functionality). However, I didn't get very far tonight - for a start, the PD height is 3" (while our beam height is 4" everywhere outside the vacuum), and I needed to put together a circuit to supply the 5V bias and +/- 15 V since the transimpedance is done on the head. I was only able to do a low-level functionality test tonight, checked that the DC voltage output varied linearly with the incident power (calibrated against an NF1611 photodiode, data will be put up later). I didn't get to measuring any noise performance - is an incandescent light bulb still shot noise limited at ~10 Hz < f < 10kHz? Some notes:

  1. The PD is DC coupled, and has a transimpedance of 1 kohm (inverting AD829 does the transimpedance).
  2. Probably a daughter board should be made that supplies the DC power voltages and rotues the output signal to something more convenient like a BNC connector. This daughter board can then also implement a DC coupled path (for monitoring) and AC coupled path (for servoing, fc to be determined).
  3. SR560 based ISS was implemented some years ago but I think the improvement was only seen above 100 Hz, and that too was marginal, the stabilized RIN was 10^-6 (monitored on an out-of-loop photodiode I think, but unsure). We'd probably want to aim for at least an order of magnitude better. Unclear at this point why more suppression wasn't possible back then, was it just insufficient loop gain, or was the sensing noise too high? To be investigated.

Unconnected to this work - this problem reared its ugly head again (i noticed it yesterday morning already actually). I don't have the energy to embark on a fix tonight, Koji is going to be in the lab all day tomorrow and so he will fix it.

ELOG V3.1.3-